# Differential Mamba

### v1

@@ 3.2 Diff-Mamba
def NormalizedDiffMamba(X, $\lambda$, Mamba):
# X.shape = (Batch, Length, Dim)
X_2 = X.repeat(1, 1, 2)
Y = Mamba(X_2)
Y_n = PostMambaNorm(Y)
Minuend = Y_N[:, :, Dim // 2:]
Subtrahend = Y_N[:, :, : Dim // 2]
O = Minuend - $\lambda$*Subtrahend
O_Prog = Out_Proj(O)
O_N = PostSubNorm(out_prog)
return O_N * (1 - lambda_init)
Figure 2: Parallel Implementation of the Normalized Differential Mamba on top of the Mamba layer.
However, as detailed in the results section, Diff S6 does not perform well and, in practice, falls short of standard Mamba layers. We suspect this arises from S6 being too simple and not functioning as a general-purpose, expressive mixing alternative to attention layers. Consequently, it fails to leverage the full potential of differential techniques. To address this limitation, we draw inspiration from Zimerman et al. [45], who demonstrated that the entire Mamba block can function as an alternative mixing mechanism to attention by formulating it as a data-controlled linear operator. In particular, the authors show that Mamba can be reformulated as implicit attention by:
$\text{Mamba}(X)=AX,\quad A=\Big{(}\text{SILU}(\text{Linear}(x))\Big{)}\hat{\alpha}\text{ diag}\Big{(}\text{Sig}(\text{Conv}(x))\Big{)})M$ (17)
where M is a matrix representing the convolution layer, and $\hat{\alpha}$ is the linear operator corresponding to S6, as formalized by Ali et al. [2].
This formulation characterizes Mamba as a data-control linear operator with a richer and more expressive implicit attention matrices. Building on this insight, we introduce Diff-Mamba, a mechanism that extends the differential approach to the full Mamba block:
[…]
Similar to Eqs. 11 and  13, this can be rewritten as a data-controlled linear operator, defined by:
[…]
@@
-$\vspace{-3pt}\mathbb{N}\text{-Diff-Mamba}(x)=\mathbb{N}(\text{Mamba}_{1}(x)-\lambda\text{Mamba}_{2}(x))$ (20)
Resulting in the normalized Diff-Mamba mechanism, which is our primary contribution.
Finally, following Ye et al. [43], we multiply the output of all variants by $1-\lambda_{\text{init}}$.
3.3 Overall Architecture
Following the Mamba architecture, our model avoids interleaving sequence mixers such as Mamba or self-attention with MLPs. Instead, it is composed entirely of Diff-Mamba blocks, augmented with standard deep learning components such as token embedding and unembedding layers, skip-connections, and normalization layers.
Efficient Implementation
Direct computation of the normalized Diff-Mamba (Eq. 20) through a naive approach can nearly double the inference latency. To mitigate this, we parallelize the computation of both the minuend and subtrahend within a single forward pass of the Mamba by partitioning the model into two independent paths. To preserve the computational complexity , parameter count, and memory footprint of the original Mamba architecture, we avoid the standard practice of channel expansion within the Mamba block, where traditionally the number of channels is doubled. Instead, we replicate the input representations across the channel dimension, effectively retaining the model size. A visualization of this parallel implementation is provided in Figure 2.
@@ 4 Experiments
Model Dataset # Layers # Params PPL (($\downarrow$( ↓)
Mamba Wikitext-103 6 167M 22.416
Diff-Mamba Wikitext-103 6 169M 22.215
@@
In this section, we empirically evaluate the effectiveness of the Diff-Mamba architecture. We begin in Section 4.1 by demonstrating that Diff-Mamba outperforms the original Mamba architecture in small-scale language modeling tasks across multiple datasets. In Section 4.2, we justify our key design decisions through a comprehensive series of ablation studies. Subsequently, in Section 4.3, we showcase the superior retrieval performance of Diff-Mamba relative to Mamba. Finally, in Section 4.4, we utilize tools from the domain of mechanistic interpretability, such as Tuned-lens [4], to examine the internal representations of Diff-Mamba in comparison to Mamba, empirically validating that our differential approach effectively reduces noise in intermediate representations. A full description of our experimental setup is provided in Appendix A.
4.1 Language Modeling
To evaluate the performance of Diff-Mamba relative to Mamba on general NLP tasks, we train both models from scratch using comparable model sizes and an identical training setup, including the same codebase [18], datasets, and hyperparameters. We focus on three widely used benchmarks: WikiText-103, Text8, and Enwik8, and experiment with models of varying depth. The final results are reported in Table 1. To provide a more comprehensive view of the optimization process, we include test curves through the training in Figure 3.
As shown in Table 1, Diff-Mamba outperforms Mamba across all evaluated environments, achieving consistently lower loss. In particular, for the 12 layer model, Diff-Mamba improves over Mamba by 0.4 perplexity on WikiText-103, 0.046 bits per byte (bpb) on Text8, and 0.041 bpb on Enwik8. For 6 layer model, Diff-Mamba improves over Mamba by 0.201 on WikiText-103, 0.02 on Text8, and 0.007 on Enwik8. Interestingly, we observe that as the number of layers increases, the differential design in Mamba becomes increasingly effective. A possible explanation for this is that, in the lower layers, the dependencies captured by the implicit attention matrices are shorter and simpler. In these cases, Mamba can manage overallocation effectively without the need for a differential mechanism. However, in the upper layers, the dependencies become more complex, spanning longer ranges [5] and exhibiting more diverse patterns. This amplifies the impact of overallocation, thereby making the benefits of the differential design more pronounced. Furthermore, the training curves in Figure 3 provide insight into the optimization properties of Diff-Mamba, showing that it consistently outperforms Mamba and achieves faster convergence. We hypothesize that this phenomenon arises from the fact that the differential design reduces the amount of noise, which appears to be critical for improving convergence [21, 44].
@@ 4.2 Ablations Analysis
4.3 Retrieval
The Diff-Transformer exhibits significantly improved retrieval capabilities compared to the original Transformer architecture. Consequently, we conduct dedicated experiments in both zero-shot and fine-tuning settings across five retrieval tasks from the BABILong benchmark [22] to evaluate whether these enhanced retrieval abilities transfer to Diff-Mamba. BABILong comprises a diverse array of reasoning tasks, including fact chaining, simple induction, and deduction, where relevant facts are embedded within lengthy natural language passages of varying context lengths. This setup provides a rigorous benchmark for evaluating models’ ability to retrieve and reason over extended contexts. In this experiment, Diff-Mamba was developed by substituting the final 12 layers of the pretrained Mamba 370M model with Diff-Mamba layers. To ensure a fair comparison, both Diff-Mamba and Mamba were fine-tuned on 2B tokens of the PG19 dataset and subsequently evaluated in a zero-shot manner on five tasks from BABILong. Furthermore, both of them were fine-tuned on BABILong tasks with up to 1k tokens to facilitate a more targeted comparison, ensuring that their learning processes were aligned with the same objectives.
Results (Figure 4) are reported on the test set exclusively. Each cell displays the ratio in which one model outperforms the other. Green indicates wins by Diff-Mamba, while red indicates wins by Mamba. As evident from the results, Diff-Mamba consistently outperforms Mamba at longer context lengths, exhibiting a slower performance degradation and a larger ratio score as context length increases. Although its performance in the zero-shot setting is partially affected by suboptimal alignment during PG19 fine-tuning, Diff-Mamba maintains strong results in the zero-shot scenario, underscoring its capacity to attend effectively to relevant contextual information. Diff-Mamba wins consistently in the BABILong finetuned test, achieving a ratio of up to 2.11, and wins the long context tasks in the zero-shot evaluation by up to 3.5. Overall, our findings demonstrate that Diff-Mamba achieves superior retrieval performance, particularly in long-context scenarios across both BABILong fine-tuned and zero-shot settings.
(a) BABILong Finetuned
(b) Zero-Shot
Figure 4: Retrieval Abilities: Comparison of Diff-Mamba and Mamba models, each with 370M parameters, under both fine-tuned and zero-shot settings across five retrieval tasks from BABILong. The x-axis represents the context length, and the y-axis corresponds to the task index. Each cell displays the ratio in which one model outperforms the other. Green cells indicate wins by Diff-Mamba, while red cells indicate wins by Mamba.
4.4 Noise Analysis in Intermediate Representations
Our empirical analysis in Section 4.3 and 4.1, as well as the results of Diff Transformer suggest that the differential design can mitigate the overallocation problem and improve general performance. To further investigate the underlying causes of this phenomenon, we analyze the model’s internal representations using tools from the field of mechanistic interpretability. In particular, we leverage the tuned lens [4] - a method designed to examine intermediate representations by training an affine probe to map activations at each layer to the model’s final prediction, thereby enabling layer-wise interpretability and insight into the model’s internal computation. Building on this tool, we measure the signal-to-noise ratio in the hidden representations of Diff-Mamba compared to standard Mamba models. The Tuned-lens tool is used to project the representations from each layer into predictions for the next token on the retrieval task. By measuring the predicted probability of the needle token, we can estimate the signal-to-noise ratio at different layers. Notably, as can be seen in Figure 5, across the majority of layers Diff-Mamba exhibits a higher signal-to-noise ratio compared to Mamba. This difference is especially pronounced in the early layers, where the predicted probability of the needle token in Diff-Mamba is several orders of magnitude higher. This analysis empirically demonstrates that Diff-Mamba produces less noisy representations, directly aligning with the key principles underlying the differential mechanism, which is designed to mitigate the overallocation problem. The linear projections for tuned-lens on both Diff-Mamba and Mamba were trained using the same configuration as in the original paper.
Figure 5: Measuring Signal-to-Noise Ratio: Noise in intermediate representations is assessed using a controlled setup integrating a well-defined retrieval task with tools from the domain of model understanding and mechanistic interpretability. The y-axis represents the probability of predicting the desired needle token, where lower values indicate higher noise. The x-axis denotes various layers within the model where intermediate noise is measured. Models are fine-tuned BABILong from Subsec. 4.3. Results show the average needle probabilities in each layer on 2k examples in test BABILong questions of up to 1k tokens.
@@ 4.5 Early Results at Medium Scale
To thoroughly evaluate Diff-Mamba, we trained both the Mamba and Diff-Mamba models from scratch, each with 370M parameters, on a 50-billion-token subset of The Pile dataset [13]. After an ablation study (Figure 10 in the Appendix), the most effective Diff-Mamba variant for scalable performance was found to be a repeated alternation of Mamba and Diff-Mamba layers throughout the network. This variant is designed to combine the strengths of both worlds: it performs better than both Mamba and Diff-Mamba, and benefits from the improved robustness, long-context, and retrieval capabilities of Diff-Mamba. Then, we evaluated the models via zero-shot tests on the LongCrawl64 dataset [7], a long-sequence subset of RedPajama-v2 [41] designed particularly for research on long-context modeling. In Figure 6 (a), the Per-token loss was calculated over the dataset following [24]. In addition, in Figure 6 (b), we calculated PPL on the test subset of The-Pile. Diff-Mamba demonstrates impressive long context capabilities, maintaining loss per token of around 9.98 across different context lengths, while Mamba increases significantly as the context grows. Furthermore, Diff-Mamba outperforms Mamba by PPL scores of 0.131, and 1.445 on The Pile and PG19 test sets correspondingly (Figure 6). We consider the positive results at medium scale promising, suggesting that alternating between Mamba and Diff-Mamba layers yields a more effective, scalable, and robust architecture.
(a) The x-axis is the token index, and the y-axis is the corresponding per-token loss.
Model Pile PG19
@@
(b) PPL results on test set of The Pile and PG19.
Figure 6: Evaluation of Mamba and Diff-Mamba in medium scale after pre-training on The Pile dataset for 50B tokens. Diff-Mamba excels in both tests.
5 Discussion: The Importance of Differential Design
@@
A key question that arises from the empirical findings presented in this paper and in Diff-Transformer is why differential design is so effective and what the underlying causes of its power are. Our work sheds light on this question by first showing that the problem of over-allocating attention to irrelevant context is not a phenomenon unique to transformers, but rather a general challenge in architectural research. Moreover, while previous work has primarily motivated the differential design through analogies to noise-canceling headphones and differential amplifiers, and supported it with strong empirical performance, we take this a step further. In Figure 5, we present a quantitative analysis using tools from the field of model understanding, showing that the intermediate representations in Diff-Mamba models exhibit a higher signal-to-noise ratio compared to their non-differential Mamba counterparts, providing empirical support for the motivation previously proposed in the literature. Yet, the underlying reasons behind the empirical success of differential design remain largely unexplored, and further investigation from both theoretical and empirical perspectives is required to advance progress in this important direction.
6 Conclusions and Limitations
@@
In this paper, we introduced Diff-Mamba, a variant of the Mamba architecture that leverages differential design principles to mitigate the problem of attention score over-allocation to irrelevant tokens, thereby enhancing overall performance. Although the experimental results show promising improvements over language modeling and retrieval tasks, we did not provide a rigorous theoretical framework explaining precisely why differential designs improve Transformer or Mamba-based LLMs. Developing such a theoretical justification is left as future work. Additionally, our results are limited to small-to-medium scale experiments due to constraints imposed by our academic budget. Finally, it remains an open question whether differential design principles can be effective in other domains, beyond NLP tasks, for instance, to domains such as computer vision, graph modeling, or time-series analysis.

### v2

@@ 3.2 Diff-Mamba
However, as detailed in the results section, Diff S6 does not perform well and, in practice, falls short of standard Mamba layers. We suspect this arises from S6 being too simple and not functioning as a general-purpose, expressive mixing alternative to attention layers. Consequently, it fails to leverage the full potential of differential techniques. To address this limitation, we draw inspiration from Zimerman et al. (2024), who demonstrated that the entire Mamba block can function as an alternative mixing mechanism to attention by formulating it as a data-controlled linear operator. In particular, the authors show that Mamba can be reformulated as implicit attention by:
$\text{Mamba}(X)=AX$ (18)
$A=\text{SILU}(\text{Linear}(x))\hat{\alpha}\text{ diag }(\text{Sig}(\text{Conv}(x)))M$
where M is a matrix representing the convolution layer, and $\hat{\alpha}$ is the linear operator corresponding to S6, as formalized by Ali et al. (2024). This formulation characterizes Mamba as a data-control linear operator with richer and more expressive implicit attention matrices. Building on this insight, we introduce Diff-Mamba, a mechanism that extends the differential approach to the full Mamba block:
$\text{Diff-Mamba}(X)={\text{Mamba}}_{1}(X)-\lambda{\text{ Mamba}}_{2}(X)$ (19)
Similar to Eqs. 12 and  14, this can be rewritten as a data-controlled linear operator, defined by:
$\text{Diff-Mamba}(X)=AX,\quad A=A_{1}-\lambda A_{2}$ (20)
@@
-$\vskip-3.0pt\mathbb{N}\text{-Diff-Mamba}(x)=\mathbb{N}(\text{Mamba}_{1}(x)-\lambda\text{Mamba}_{2}(x))$ (21)
Resulting in the normalized Diff-Mamba mechanism, which is our primary contribution. Finally, following Ye et al. (2024), we multiply the output of all variants by $1-\lambda_{\text{init}}$.
@@ 4 Experiments
In this section, we empirically evaluate the effectiveness of the Diff-Mamba architecture. We begin in Section 4.1 by demonstrating that Diff-Mamba outperforms the original Mamba architecture in small-scale language modeling tasks across multiple datasets. In Section 4.2, we justify our key design decisions through a comprehensive series of ablation studies. Subsequently, in Section 4.3, we use carefully designed synthetic tasks that are predictive of model behavior in large scale settings. Then, in Section 4.4 we train the models at medium scale and test long-context capabilities. Afterwards, in Section 4.5 we showcase the superior retrieval performance of Diff-Mamba relative to Mamba. Finally, in Section 4.6, we utilize tools from the domain of mechanistic interpretability, such as Tuned-lens (Belrose et al., 2023), to examine the internal representations of Diff-Mamba in comparison to Mamba, empirically validating that our differential approach effectively reduces noise in intermediate representations. A full description of our experimental setup, as well as an efficiency analysis, is provided in Appendix A and Appendix B.
Model Dataset # Layers # Params PPL $(\downarrow$)
Mamba Wikitext-103 6 167M 22.357
Diff-Mamba Wikitext-103 6 169M 22.282
@@
4.1 Language Modeling
To evaluate the performance of Diff-Mamba relative to Mamba on general NLP tasks, we train both models from scratch using comparable model sizes and an identical training setup, including the same codebase (Gu et al., 2022), datasets, and hyperparameters. We focus on three widely used benchmarks: WikiText-103, Text8, and Enwik8, and experiment with models of varying depth. The final results are reported in Table 1. To provide a more comprehensive view of the optimization process, we include test curves through the training in Figure 2.
As shown in Table 1, Diff-Mamba outperforms Mamba across all evaluated environments, achieving consistently lower loss. In particular, for the 12 layer model, Diff-Mamba improves over Mamba by 0.4 perplexity on WikiText-103, 0.046 bits per byte (bpb) on Text8, and 0.041 bpb on Enwik8. For 6 layer model, Diff-Mamba improves over Mamba by 0.075 on WikiText-103, 0.02 on Text8, and 0.007 on Enwik8. Interestingly, we observe that as the number of layers increases, the differential design in Mamba becomes increasingly effective. A possible explanation for this is that, in the lower layers, the dependencies captured by the implicit attention matrices are shorter and simpler. In these cases, Mamba can manage overallocation effectively without the need for a differential mechanism. However, in the upper layers, the dependencies become more complex, spanning longer ranges (Ben-Kish et al., 2025a) and exhibiting more diverse patterns. This amplifies the impact of overallocation, thereby making the benefits of the differential design more pronounced. Furthermore, the training curves in Figure 2 provide insight into the optimization properties of Diff-Mamba, showing that it consistently outperforms Mamba and achieves faster convergence. We hypothesize that this phenomenon arises from the fact that the differential design reduces the amount of noise, which appears to be critical for improving convergence (Johnson and Zhang, 2013; Zhang et al., 2019).
@@ 4.2 Ablations Analysis
4.3 Unit Tests as Predictors of Scaling
Training LLMs at scale is prohibitively resource-intensive. Nevertheless, prior work has established that carefully designed small-scale capability evaluations Gupta et al. (2022) are predictive of model behavior in large-scale Poli et al. (2024). To this end, we follow the MAD pipeline Poli et al. (2024) and adopt a rigorous suite of synthetic token manipulation tasks, that serve as capability unit tests. These tasks allow us to isolate and assess core mechanisms that underpin scalable performance. We investigate Mamba and Diff-Mamba. Diff-Mamba model consistently demonstrates superior results (Figure 5), providing strong evidence that the architecture is well-positioned to retain and extend these capabilities when trained at larger scales. Both models achieve the highest accuracy on in-context-recall (ICR), noisy ICR, and selective-copying tasks, while Diff-Mamba outperforms Mamba on the rest of the tasks, specifically up to 10.6% improvement in compression, 80% in fuzzy ICR, and 0.2% in memorization. Diff-Mamba surpasses Mamba by 3.7% in total.
@@ 4.4 Early Results at Medium Scale
To thoroughly evaluate Diff-Mamba, we trained both Mamba and Diff-Mamba from scratch, each with 370M parameters, on a 50B token subset of The Pile dataset (Gao et al., 2020). After an ablation study (Figure 7 in the Appendix), the most effective Diff-Mamba variant for scalable performance was found to be a repeated alternation of Mamba and Diff-Mamba layers throughout the network. This hybrid variant performs better than both models with fully Mamba or Diff-Mamba layers.
Next, we evaluated the models via zero-shot tests on the LongCrawl64 dataset (Buckman, 2024), a long-sequence subset of RedPajama-v2 (Weber et al., 2024) designed particularly for research on long-context. In Figure 3 (a), per-token loss was calculated over the dataset following (Lin et al., 2025). In addition, in Figure 3 (b), we calculated PPL on the test subset of The Pile and PG19. Diff-Mamba demonstrates impressive long context capabilities, maintaining per-token loss of around 9.98 across different context lengths, while Mamba increases significantly as the context grows. Furthermore, Diff-Mamba outperforms Mamba by PPL scores of 0.131, and 1.445 on The Pile and PG19 test sets respectively. We consider the positive results at medium scale promising, suggesting that alternating between Mamba and Diff-Mamba layers yields a more effective, scalable, and robust architecture.
Model The Pile $(\downarrow$) PG19 $(\downarrow$)
@@
Figure 3: Diff-Mamba excels in both tests. (a) The x-axis is the token index, and the y-axis is the corresponding per-token loss. (b) PPL results on test set of The Pile and PG19.
Figure 4: Retrieval Abilities: Comparison of Diff-Mamba and Mamba models across five retrieval tasks from BABILong. X-axis represents the context length, and y-axis corresponds to the task index. Each cell displays the ratio in which one model outperforms the other. Green cells indicate wins by Diff-Mamba, while red cells indicate wins by Mamba.
Figure 5: Performance comparison on synthetic token manipulation tasks. We evaluate Mamba and Diff-Mamba architectures across six synthetic capability benchmarks. Diff-Mamba consistently outperforms Mamba, with notable improvements in Fuzzy ICR of 80.0% and Compression with 10.6%. Stars indicate the best model per task.
4.5 Retrieval
The Diff-Transformer exhibits significantly improved retrieval capabilities compared to the original Transformer. Consequently, we conduct experiments across five retrieval tasks from the BABILong benchmark (Kuratov et al., 2024) to evaluate whether these enhanced abilities transfer to Diff-Mamba. BABILong comprises a diverse array of reasoning tasks, including fact chaining, simple induction, and deduction, where relevant facts are embedded within lengthy natural language passages of varying context lengths. This setup provides a rigorous benchmark for evaluating models’ ability to retrieve and reason over extended contexts. To ensure a fair comparison, both Diff-Mamba and Mamba (from Section 4.4) were fine-tuned on BABILong tasks with up to 1k tokens to facilitate a more targeted comparison, ensuring that their learning processes were aligned with the same objectives.
Results (Figure 4) are reported on the test set. Each cell displays the ratio in which one model outperforms the other. Green indicates wins by Diff-Mamba, while red indicates wins by Mamba. As evident from the results, Diff-Mamba outperforms Mamba, exhibiting a slower performance degradation and a larger ratio score as context length increases. Diff-Mamba wins consistently, achieving a ratio of up to 3.95. Our findings demonstrate that Diff-Mamba achieves superior retrieval performance, particularly in long-context scenarios.
Figure 6: Measuring Signal-to-Noise Ratio: The y-axis represents the probability of predicting the desired needle token, where lower values indicate higher noise. The x-axis denotes various layers within the model where intermediate noise is measured. Results show the average needle probabilities in each layer on 1k examples in BABILong questions of 1k-2k tokens.
4.6 Noise in Intermediate Representations
Our empirical analysis in previous sections, as well as the results of Diff-Transformer suggest that the differential design can mitigate the overallocation problem and improve general performance. To further investigate the underlying causes of this phenomenon, we analyze the model’s internal representations using tools from the field of mechanistic interpretability. In particular, we leverage Tuned-lens (Belrose et al., 2023) - a method designed to examine intermediate representations by training an affine probe to map activations at each layer to the model’s final prediction, thereby enabling layer-wise interpretability and insight into the model’s internal computation. Building on this tool, we measure the signal-to-noise ratio in the hidden representations of Diff-Mamba compared to standard Mamba models. The Tuned-lens tool projects logits from each layer into predictions for the next token on the retrieval task. By measuring the predicted probability of the needle token, we can estimate the signal-to-noise ratio at different layers. Notably, as can be seen in Figure 6, across the majority of layers Diff-Mamba exhibits a higher signal-to-noise ratio compared to Mamba. This difference is especially pronounced in the early to mid layers, where the predicted probability of the needle token in Diff-Mamba is substantially higher. This analysis empirically demonstrates that Diff-Mamba produces less noisy representations, directly aligning with the key principles underlying the differential mechanism, which is designed to mitigate the overallocation problem.
5 Discussion: Why Differential Design
@@
A key question that arises from the empirical findings presented in this paper and in Diff-Transformer is why differential design is so effective and what the underlying causes of its power are. Our work sheds light on this question by first showing that the problem of over-allocating attention to irrelevant context is not a phenomenon unique to transformers, but rather a general challenge in architectural research. Moreover, while previous work has primarily motivated the differential design through analogies to noise-canceling headphones and differential amplifiers, and supported it with strong empirical performance, we take this a step further. In Figure 6, we present a quantitative analysis using tools from the field of model understanding, showing that the intermediate representations in Diff-Mamba exhibit a higher signal-to-noise ratio compared to their non-differential Mamba counterparts, providing empirical support for the motivation previously proposed in the literature. Yet, the underlying reasons behind the empirical success of differential design remain largely unexplored, and further investigation from both theoretical and empirical perspectives is required to advance progress in this important direction.
6 Conclusions
In this paper, we introduced Diff-Mamba, a variant of the Mamba architecture that leverages differential design principles to mitigate the problem of attention score over-allocation to irrelevant tokens, thereby enhancing overall performance, with a particular emphasis on long-context, retrieval and robustness capabilities. Diff-Mamba advances architectural research on Mamba variants by introducing an inductive bias that enhances the signal-to-noise ratio, a property that may be critical at scale, as hallucinations remain one of the most significant challenges in LLMs.
7 Limitations
@@
Although the experimental results show promising improvements over language modeling and retrieval tasks, we did not provide a rigorous theoretical framework explaining precisely why differential designs improve Transformer or Mamba-based LLMs. Developing such a theoretical justification is left as future work. Additionally, our results are limited to small-to-medium scale experiments due to constraints imposed by our academic budget. Finally, it remains an open question whether differential design principles can be effective in other domains, beyond NLP tasks, for instance, to domains such as computer vision, graph modeling, or time-series analysis.
Ethics Statement
Our work aims to improve the robustness of Mamba models by increasing the signal-to-noise ratio through differential design. In the broader context, this contribution supports the development of more efficient, reliable, and trustworthy LLMs. Additionally, our findings highlight when and how differential design is effective and offer practical guidance for incorporating inductive biases that lead to less noisy architectures. Thus, we conclude that our approach contributes to the principled design of safer and more reliable LLMs.